### v2

# Securing Smart Contract Languages with a Unified Agentic Framework for Vulnerability Repair in Solidity and Move

Abstract
The rapid growth of the blockchain ecosystem and the increasing value locked in smart contracts necessitate robust security measures. While languages like Solidity and Move aim to improve smart contract security, vulnerabilities persist. This paper presents Smartify, a novel multi-agent framework leveraging Large Language Models (LLMs) to automatically detect and repair vulnerabilities in Solidity and Move smart contracts. Unlike traditional methods that rely solely on vast pre-training datasets, Smartify employs a team of specialized agents working on different specially fine-tuned LLMs to analyze code based on the underlying programming concepts and language-specific security principles. We evaluated Smartify on a dataset for Solidity and a curated dataset for Move, demonstrating its effectiveness in fixing a wide range of vulnerabilities. Our experimental results show that Smartify (Gemma2+Codegemma) achieves state-of-the-art performance, surpassing existing LLMs and even enhancing the capabilities of general-purpose models, such as Llama 3.1. Notably, Smartify can incorporate language-specific knowledge, such as the nuances of Move, without requiring massive language-specific pretraining datasets. This work offers a detailed analysis of the performance of various LLMs on smart contract repair, highlighting the strengths of our multi-agent approach and providing a blueprint for developing more secure and reliable decentralized applications in the growing blockchain landscape. We also provide a detailed description to extend the proposed technology to other similar use cases.
1 Introduction
Smart contracts, self-executing agreements with terms directly written into code, have emerged as a cornerstone of blockchain technology nath2014web; ray2023web3. Their ability to automate transactions and eliminate intermediaries has led to widespread adoption in various sectors, including finance, supply chain management, and healthcare zheng2018blockchain; karanjai2021conditional; kaleem2021event. However, the increasing complexity of smart contracts has given rise to a growing concern: security vulnerabilities vacca2021systematic. These vulnerabilities, often stemming from coding errors or design flaws, can be exploited by malicious actors, leading to significant financial losses and damage to the reputation of blockchain projects.
The financial implications of smart contract vulnerabilities are substantial. Reports indicate that cumulative losses from attacks against Ethereum smart contracts alone have exceeded USD 3.1 billion by 2023 li2023smart. In the DeFi space, an estimated \$9.04 billion has been stolen due to vulnerabilities wronka2023financial. Notable incidents like the DAO hack of 2016, resulting in a \$55 million loss popper2016hacking, and the Poly Network hack in 2021, where over \$600 million was stolen polyhack, underscore the critical need for robust security measures.
Traditional security auditing methods, although essential, often face limitations in terms of accuracy and scalability. This has spurred the exploration of automated techniques for vulnerability detection 10.1145/3238147.3238177; wang2020contractwardand repair, with Large Language Models (LLMs) emerging as a promising solution joshi2023repair. LLMs trained on vast datasets of code can learn to understand and generate code that adheres to specific programming paradigms and best practices. However, most of the tools available for smart contract safety are language-specific, especially Solidity dannen2017solidity. For other languages, existing tools often require scanning of compiled bytecode song2024empirical.
Apart from Solidity, Move blackshear2019move has gained significant traction due to its strong focus on security. Its cutting-edge features, including a custom data type for secure operations, robust access controls via Move modules, and unique memory safety features blackshear2022move, have been particularly noteworthy. Moreover, the Move Prover, a native security framework, provides an additional layer of protection dill2022fast. Notably, several prominent blockchain platforms, such as Starcoin starcoin, Aptos devaptos, and Sui blackshear2024sui, have already adopted Move.
However, despite its promising architecture, the real-world security performance of Move modules remains largely untested. Unlike Solidity-based smart contracts, which have been extensively studied through empirical research and surveys, there is a scarcity of research focused specifically on Move modules. Although some methodologies have been proposed for identifying defects in Move modules or conducting formal verification keilty2022model; park2024securing, and empirical analysis song2024empirical, a significant knowledge gap persists. Specifically, large-scale investigations into the frequency of defects in real-world Move modules and identifying and repairing potential vulnerabilities are lacking, highlighting the need for further research in this area.
This paper introduces Smartify, a framework that moves beyond naive LLM application by proposing a novel paradigm for structured, automated repair. We introduce Smartify not merely as a "multi-agent framework" but as a structured, collaborative system that mimics the workflow of an expert human security audit team. This contrasts with existing LLM-powered tools like ContractTinker, which utilize a linear Chain-of-Thought (CoT) reasoning process to break down the repair task. Smartify’s novelty lies in its process-centric design, where specialized agents (Auditor, Architect, Code Generator, Refiner, Validator) assume distinct roles within a delegative and iterative workflow. The Architect agent, for instance, does not simply reason about the next step; it formulates a comprehensive repair strategy that is then executed by other agents. Crucially, the Refiner-Validator duo establishes a feedback loop for iterative quality assurance, a feature essential for generating trustworthy patches.
In this paper, our aim is to answer the following research questions related to software engineering using AI agents and the landscape of complex smart contract reasoning.
  • RQ1: Do the present state-of-the-art LLMs can explain a Smart Contract code correctly?
  • RQ2: Can they detect and explain bad coding practices or specific mistakes leading to bugs or vulnerabilities in a smart contract code?
  • RQ3: Can we encode programming language-specific knowledge to train the LLMs to understand unsafe and buggy codes in detail enough to repair them?
  • RQ4: Can LLMs repair the bugs and fix the vulnerability?
  • RQ5: Does the proposed post-training framework be generalizable to a larger set of pre-trained LLMs?
The main contributions of our work are the following:
  • We introduce a novel role-based multi-agent architecture for smart contract repair that models the collaborative workflow of a security audit team, enhancing structured reasoning over monolithic LLM approaches.
  • We propose a rigorous methodology combining specialized fine-tuned models for deep vulnerability analysis, Retrieval-Augmented Generation (RAG) for language-specific context, and an iterative self-refinement loop for ensuring patch quality.
  • We conduct the first multi-faceted empirical evaluation of an LLM-based repair tool for both Solidity and Move, using metrics that assess not only correctness but also exploit mitigation effectiveness, semantic preservation, and code quality.
  • We present a comprehensive ablation study that systematically dissects our framework to quantify the distinct contribution of each architectural component to the overall repair performance.
2 Related Work
This section provides a critical review of the existing landscape in smart contract security, positioning Smartify relative to traditional analysis techniques and emerging LLM-based repair methodologies.
2.1 Traditional Smart Contract Security Analysis
Automated security analysis for smart contracts has traditionally been dominated by static, dynamic, and formal verification techniques. Each approach offers distinct advantages but also possesses inherent limitations, particularly when confronted with complex, logic-based vulnerabilities.
2.2 Smart Contract Security Auditing
Various tools and techniques have been developed for detecting vulnerabilities in smart contracts:
Static Analysis Tools: Tools like Mythril muellerfile and Slither feist2019slither analyze contract source code to identify potential vulnerabilities. They perform symbolic execution and taint analysis to detect patterns associated with common vulnerabilities.
Dynamic Analysis Tools: Tools like Manticore mossberg2019manticore and Echidna grieco2020echidna execute contracts with various inputs to uncover runtime errors. They use fuzzing and symbolic execution techniques to explore different execution paths and identify potential issues.
Formal Verification: This approach uses mathematical techniques to rigorously prove the correctness of a contract’s code against a formal specification. Tools like KEVM hildenbrandt2018kevm and CertiK’s DeepSEA have been developed for formal verification of smart contracts zhong2020move.
While these tools are valuable, they often have limitations in accuracy, scalability, and the ability to handle the complexities of real-world smart contracts.
2.3 LLM-Powered Smart Contract Repair: A Taxonomy
The application of LLMs to automated program repair is a rapidly advancing field, with several distinct methodologies emerging.
2.3.1 Monolithic LLM Approaches
Early efforts in this domain involved using large, general-purpose LLMs like GPT-3 with intricate, zero-shot or few-shot prompts to generate patches. While demonstrating feasibility, these approaches often lack the domain-specific knowledge required for the high-stakes environment of smart contracts. They are prone to generating syntactically correct but semantically flawed or insecure code, as they lack a deep, ingrained understanding of blockchain-specific security paradigms blackshear2022move; bobadilla2025automated.
2.3.2 Chain-of-Thought (CoT) and Static Analysis Integration
To address the limitations of monolithic models, more structured approaches have been developed. A prominent example is ContractTinker, a tool designed for real-world smart contract repair wang2024contracttinker. ContractTinker employs a Chain-of-Thought (CoT) mechanism to guide an LLM through a sequence of reasoning steps: vulnerability localization, analysis, and patch generation. To ground the LLM’s reasoning and mitigate hallucination, it integrates static analysis techniques, including dependency analysis and program slicing, to provide relevant context from audit reports and the source code itself.
While ContractTinker represents a significant advancement by imposing a logical structure on the repair process, its CoT approach remains a fundamentally linear and sequential reasoning pipeline. In contrast, Smartify’s architecture is delegative and iterative. The Architect agent formulates a comprehensive, high-level plan, which is then delegated to specialized agents for execution. Furthermore, the explicit Refiner-Validator loop introduces a crucial mechanism for feedback and iterative quality improvement, a feature not explicitly detailed in the ContractTinker workflow. This architectural distinction moves beyond a simple chain of thought to a collaborative problem-solving process.
2.3.3 Multi-Agent Systems for Code Tasks
The concept of using multiple LLM-based agents to collaborate on complex tasks has gained traction in the broader software engineering domain, with applications in areas like code translation and generation huang2024opencoder; jiang2018contractfuzzer. These systems leverage the principle of specialization, assigning different roles or sub-tasks to individual agents to achieve a more robust and accurate outcome than a single agent could.
While Smartify aligns with this general trend, its novelty lies in its domain-specific agent roles tailored explicitly for the vulnerability repair workflow. Instead of a generic "translator" or "coder" agent, Smartify’s agents embody the distinct functions of a human security audit team: the Auditor for analysis, the Architect for strategic planning, the Code Generator for implementation, and the Refiner/Validator for quality assurance. This specialization allows for a more nuanced and effective approach to the highly specific and critical task of securing smart contracts.
Our proposed framework, Smartify, addresses these challenges by combining the strengths of specialized LLMs within a multi-agent architecture. It leverages language-specific fine-tuning, safety classifiers, and Retrieval-Augmented Generation (RAG) to enhance the accuracy and security of generated code repairs.
In the following sections, we detail the architecture of Smartify, describe the experimental setup, present the evaluation results, and discuss the implications of our findings for the future of smart contract security.
3 Data Collection and Analysis Methodology
This research employs a multi-faceted approach to investigate the security of smart contracts, focusing on both Solidity and Move programming languages. The methodology encompasses collecting and analyzing two distinct datasets: Solidity-based, Move-based source code. Each dataset serves a specific purpose in addressing the research questions and contributing to a comprehensive understanding of smart contract vulnerabilities.
3.1 Importance of Dataset Categorization
For several reasons, categorizing the datasets based on programming language (Solidity and Move) and code representation is crucial. It allows for a focused analysis of language-specific vulnerabilities and coding practices. As a more mature language, Solidity exhibits a different vulnerability landscape than the newer Move language. Examining them separately enables the identification of unique challenges and security considerations associated with each language.
3.2 Dataset Descriptions
3.2.1 Solidity-based Dataset
This dataset comprises a collection of vulnerable Solidity smart contracts sourced from the "Not-So-Smart Contracts" repository curated by Trail of Bits githubGitHubCryticnotsosmartcontracts. This repository is renowned for its comprehensive set of contracts that intentionally exhibit a variety of common vulnerabilities. These vulnerabilities were chosen for inclusion because of their prevalence in real-world decentralized applications and their representation of typical errors during smart contract development. The dataset contains 60 vulnerable contracts, encompassing 8 distinct vulnerability categories. Table 1 shows these categories’ distribution.
Table 1: Distribution of Vulnerabilities in the Solidity Dataset.
Vulnerability Type Number of Contracts Percentage (%)
Reentrancy $15$ $25.0$
Integer Overflow/Underflow $10$ $16.7$
Denial of Service (DoS) 88 $13.3$
Access Control Issues $12$ $20.0$
Uninitialized Storage Pointers 55 $8.3$
Tx.origin Misuse 33 $5.0$
Timestamp Dependency 44 $6.7$
Gas Limit and Out-of-Gas Vulnerabilities 33 $5.0$
Total 60 100.0
3.2.2 Move-Based Dataset (Source Code)
This dataset encompasses the source code of 92 real-world Move projects, comprising 652 individual modules. These projects were part of Aptos devaptos, Sui blackshear2024sui, and Starcoin starcoin. These projects span various application domains, as depicted in Table 2. The total number of Move projects is 92, and the total number of Move modules within these projects is 652.
Table 2: Distribution of Move Projects by Application Domain.
Application Domain Number of Projects Percentage (%)
Decentralized Finance $41$ $44.6$
Token $22$ $23.9$
Bridge $18$ $19.6$
Library 33 $3.3$
Infrastructure 33 $3.3$
Other 55 $5.4$
Total 92 100.0
For both the Move-based datasets, we utilize Song et al.’s song2024empirical work to compare the vulnerability detection part. While the prior work is directed towards detection, the same dataset helps us compare Smartify’s performance on both detection and repair.
3.3 Evaluation of Smartify
Smartify is designed with two core functionalities: detecting and repairing unsafe coding patterns in smart contracts. We utilize the previously described datasets to evaluate these capabilities rigorously, encompassing both Solidity and Move code. The evaluation process focuses on the complete output of Smartify rather than individual components, reflecting its nature as an integrated solution for smart contract security. Performance is measured using the Pass@1 score.
3.4 Agent-Based Code Repair Process for Smart Contracts
Our approach leverages a multi-agent system inspired by established software development methodologies but tailored explicitly for the automated repair of Solidity and Move smart contracts. This system employs five specialized agents: an Auditor, an Architect, a Code Generator, a Refiner, and a Validator. The process incorporates a self-refinement loop and a final validation step, ensuring high accuracy and security. Each agent plays a distinct role in a structured workflow, detailed below.
Figure 1: Agentic architecture of Smartify.
  • Auditor: This agent is the cornerstone of security analysis. It is fine-tuned on a comprehensive corpus of Solidity and Move code documentation, encompassing syntax, semantics, and best practices. Furthermore, it is safety-aligned using a classifier adapted from Google’s Responsible AI toolkit. This classifier has been meticulously modified to enforce language-specific rules and safe coding practices, effectively preventing the generation of unsafe or unsupported code constructs.
    This alignment is of paramount importance. For Move, it ensures that the generated code strictly adheres to the conventions of the target blockchain (e.g., Sui or Aptos). It prevents the accidental introduction of elements from one Move variant into another or the inclusion of unsupported Rust paradigms. This is crucial because Move, derived from Rust, has unique features and limitations. For Solidity, it enforces established security best practices and prevents the generation of code patterns known to be vulnerable.
    The Auditor’s primary responsibility is to meticulously scan the input smart contract code (either Solidity or Move) to identify potential vulnerabilities and unsafe patterns. Its secondary, yet vital, role is to serve as the final Validator of the repaired code.
  • Architect: This agent receives the output from the Auditor, which includes a detailed report of identified vulnerabilities and unsafe code segments. The Architect’s role is to devise a high-level strategic plan for addressing these issues. This plan does not involve generating code directly. Instead, it outlines the necessary modifications, refactoring, and improvements to rectify the identified problems. This plan is a comprehensive blueprint for the Code Generator, guiding the code repair process.
  • Code Generator: This agent is a general-purpose code LLM. Its strength lies in its ability to leverage Retrieval-Augmented Generation (RAG) from two distinct data stores, one dedicated to Solidity and the other to Move. These data stores contain a collection of best practices and relevant documentation for the respective programming languages.
    Using the Architect’s plan as a guide, the Code Generator selects and adapts relevant examples from the appropriate RAG datastore. This dynamic, context-aware retrieval of few-shot examples significantly enhances the Code Generator’s ability to produce accurate and secure code repairs. It ensures the generated code adheres to language-specific conventions and incorporates established best practices.
  • Refiner: This agent’s role is to enhance the code quality produced by the Code Generator. It achieves this through iterative self-refinement, essentially acting as its critic. The Refiner uses the same underlying LLM as the Code Generator but with a different prompt that focuses on improving the code quality based on best practices and potential improvements that it might detect from a higher level.
  • Validator: This agent acts as a final checkpoint. After the refinement stage, it re-employs the Auditor agent to re-evaluate the code. The Validator’s objective is to ensure that all the previously identified vulnerabilities have been adequately addressed and that no new vulnerabilities have been introduced during the repair and refinement process.
4 Smartify System Architecture and Workflow
Smartify operates through a five-agent system designed for automated intelligent contract vulnerability detection and repair. The system functions as shown in Figure 1.
The Smartify system operates in a five-phase process to automatically repair smart contract code. Firstly, in the Input & Initial Audit phase, the smart contract code, written in either Solidity or Move, is fed into the system. The Auditor, an LLM based on Gemma2 9B, analyzes the code to detect potential vulnerabilities and produces a report detailing its findings. Secondly, during Repair Planning, the Architect receives this vulnerability report and formulates a high-level repair plan that outlines the necessary code modifications to address the identified issues. Thirdly, in Code Generation & Refinement, an LLM called CodeGemma, which has been fine-tuned for code generation and is equipped with Retrieval-Augmented Generation (RAG) capabilities, takes the lead. It utilizes separate Move RAG and Solidity RAG components to provide language-specific context. The Code Generator, part of CodeGemma, uses the repair plan to generate the modified code, selecting the appropriate RAG based on the input language and able to perform Solidity to Move translation when necessary.
Subsequently, a Self-refinement process is initiated, and the Refiner component iteratively improves the generated code’s quality, readability, and efficiency. Fourthly, in the Validation phase, the Validator (the same agent as the Auditor) performs a final security audit on the refined code to ensure all identified vulnerabilities have been resolved. Finally, the system outputs the repaired smart contract code.
The process may iterate to step 3 or 4 if the Validator identifies any issues. Each step is vital in ensuring the smart contract code’s accurate and secure repair. The workflow is designed to be efficient and effective, leveraging each agent’s strengths to achieve the desired outcome.
4.1 Agent Prompting Strategy
The agents within Smartify are driven by carefully crafted prompts that guide their actions and ensure consistent performance. We employ a standardized prompt template adapted from established practices in LLM-based agent systems. The template is structured as follows:
Prompt Template Role: You are a [role] specializing in [Solidity/Move] smart contracts. Task: [task] Instruction: Based on the provided Context, please follow these steps: [numbered steps] Context:
This template is broken down into the following components.
Each agent in our framework is defined by four key components: the Role, which designates the agent’s specific function (such as Auditor, Architect, or Code Generator); the Task, which outlines the agent’s particular objectives; the Instruction, which provides detailed step-by-step guidance using chain-of-thought reasoning; and the Context, which encompasses all necessary information including input code, audit reports, architectural plans, RAG datastore examples, and inter-agent conversation history.
Table 3 shows how this template is adapted for each agent.
Table 3: Agent Prompts for Smart Contract Repair.
Role Task Instruction Context
Auditor Identify vulnerabilities and unsafe patterns in Solidity/Move code. Analyze the code for security vulnerabilities and generate a detailed report. Input smart contract code (Solidity/Move).
Architect Create a high-level plan to address vulnerabilities identified by the Auditor. Review the Auditor’s report and develop a plan outlining necessary modifications. Auditor’s report.
Code Generator Generate Repaired Solidity/Move code based on the Architect’s plan and RAG examples. Consult the Architect’s plan, retrieve examples from the RAG datastore, and generate repaired code. Architect’s plan, Solidity/Move code examples from RAG.
Refiner Iteratively refine the generated code to improve quality and efficiency. Review the generated code, identify areas for improvement, and refine accordingly. Generated code, previous iteration code (if any).
Validator Perform a final security check on the repaired code. Analyze the repaired code for vulnerabilities, verify issue resolution, and ensure no new vulnerabilities. Repaired smart contract code.
4.2 Hardware and Model Fine-tuning
The development and deployment of Smartify leveraged a heterogeneous computing environment, utilizing high-performance GPUs for computationally intensive tasks and a more resource-efficient setup for inference.
4.2.1 Fine-tuning Setup
  • Hardware: Fine-tuning leveraged a cluster of four NVIDIA A100 GPUs for computationally demanding pattern learning in Solidity and Move code.
  • Model: Based on the Gemma 9B model, selected for strong code-related task performance and fine-tuning adaptability, particularly in instruction following. Fine-tuned on a dataset of Solidity and Move code, vulnerability examples, best practices, and documentation, augmented with outputs from earlier pipeline stages to enhance safety issue detection.
  • Training Recipe: Supervised learning paradigm. Trained to predict correct outputs (e.g., vulnerability reports, safe code patterns) from inputs (e.g., Solidity/Move code, vulnerability descriptions).
    • Data Pre-processing: Tokenization, normalization, and input-output pair creation ensured data consistency and quality.
    • Hyperparameter Optimization: Learning rate (1e-5), batch size (8, due to memory constraints), and training epochs (5, as validation loss plateaued) optimized via grid search and manual tuning.
    • Regularization: Dropout and weight decay used to prevent overfitting and improve generalization.
    • Evaluation Metrics: Accuracy, precision, recall, and F1-score on a held-out validation set monitored model performance.
4.2.2 Inference Setup
  • Hardware: Inference was performed on a single NVIDIA RTX 4090 GPU, balancing performance and cost-effectiveness for real-time code repair.
  • Models:
    • Code Generator and Refiner: These agents utilize a fine-tuned CodeGemma model, initially pre-trained on a limited Move corpus and further instruction-tuned to follow Architect-generated "recipe" patterns. Fine-tuning on Architect outputs ensured it understood these instructions, and pre-training on a limited Move corpus ensured basic syntax understanding.
    • Comparison Model: A stock Llama 3.1 model was used in some experiments for comparative analysis, helping assess the gains from fine-tuning and instruction tuning.
4.2.3 Key Considerations
  • A balance between performance requirements, resource availability, and cost considerations drove the choice of hardware and models.
  • The fine-tuning process for the Auditor was particularly resource-intensive due to the complexity of the task and the size of the model.
  • The use of a smaller, more efficient GPU for inference makes the system more accessible for practical deployment.
  • The comparison with a stock Llama 3 model provides valuable insights into the effectiveness of our fine-tuning and instruction-tuning strategies.
This heterogeneous setup, combining high-performance GPUs for training and a more efficient GPU for inference, allows Smartify to effectively address the computational demands of both model development and deployment. The detailed description of the fine-tuning process provides transparency and allows for replication of our results.
5 Experimental Results and Discussion
We run our experiments as defined in Section 3.3. We report the results as well as the empirical performance of our models. Through that, we will try to answer our Research Questions individually in this section.
Along with Smartify, we have run the benchmark for the following models.
Table 4: Comparison of Code and Non-Code Models.
Model Name Parameters Quantization Code Model
Granite-Code 8B FP16 Yes
CodeGemma 7B FP16 Yes
DeepSeek-Coder-v2 N/A N/A Yes
StarCoder2 15B FP16 Yes
CodeGeex4 13B N/A Yes
CodeStral 7B FP16 Yes
DeepSeek-Coder 33B N/A Yes
CodeLlama roziere2023code 13B N/A Yes
CodeQwen 7B Q8_0 Yes
Qwen2.5-coder 2.5B N/A Yes
Gemma2 N/A N/A Yes
Gemma2:27b 27B FP16 Yes
Llama3.2 3.2B FP16 No
OpenCoder 8B FP16 Yes
Llama3.3 3.3B FP16 No
The models were chosen according to the top 8 models at Hugging Face Big Code Leaderboard huggingfaceCodeModels at the time of this work, and also adding general-purpose models, which are supposed to be better at reasoning.
5.1 Solidity
This section presents the evaluation results of various code generation models on repairing vulnerabilities in Solidity smart contracts, specifically focusing on the “Not So Smart Contracts” dataset from the Trail of Bits GitHub repository. This dataset is a collection of intentionally vulnerable Solidity contracts designed to test the ability of automated tools to detect and repair common security flaws. It contains diverse vulnerabilities, including re-entrancy, integer overflow/underflow, access control issues, and timestamp dependence, among others. The dataset has been publicly available for a significant period, raising the possibility that some or all of its contents might be present in the pre-training data of the evaluated models. We analyze the performance of these models based on two key metrics: the number of vulnerabilities fixed and the average inference time, as summarized in Table 5 and Figure 2. We also introduce our framework, Smartify, and demonstrate its effectiveness in enhancing model performance.
Figure 2: Code Repair: Solidity.
Table 5: Performance of Code Generation Models on Vulnerability Repair.
Model Name Vuln. Fixed Avg. Time (s)
CodeGeex-4 11 95.50
CodeGemma 16 96.50
CodeLlama 3 243.93
CodeQwen 5 141.05
CodeStral 13 295.23
DeepSeekCoder-33b 1 411.75
DeepSeek-V2 9 19.42
Gemma2-9b 13 108.30
Gemma2-27b 14 304.27
Granite-Code 14 90.37
Llama3.2 1 37.09
Llama3.3-70b 13 741.10
OpenCoder 1 94
Qwen-2.5-Code 13 79.72
StarCoder2 0 89.10
Smartify (Gemma2+CodeGemma) 16 112.30
Smartify (Gemma2+Llama3.1) 14 267.80
The results reveal significant performance disparities among the evaluated models. Among the pre-trained models for Solidity CodeGemma surprisingly emerges as a top performer, successfully fixing 16 vulnerabilities with a relatively low average inference time of 96.5 seconds. This suggests that CodeGemma possesses a strong ability to understand and rectify code vulnerabilities while maintaining reasonable efficiency. However since most of these Solidity smart contracts were part of open Githubs repositories, there can be a strong possibility fo these already being part of the pertaining data. Our proposed framework, Smartify (Gemma2+CodeGemma), achieves comparable performance, also fixing 16 vulnerabilities, albeit with a slightly higher average inference time of 112.3 seconds. This increased time is likely due to its iterative multi-agent process, which enables Smartify to leverage the complementary strengths of Gemma2 and CodeGemma, resulting in robust and reliable fixes.
While Smartify does not immediately show any benefits over CodeGemma here, we can notice that the same Smartify framework when applied to Llama3.1 without any fine-tuning (unlike the Smartify with CodeGemma) still gives considerable performance boost over vanilla.
Figure 3: Move Code Repair.
Conversely, models like CodeLlama, CodeQWen, DeepSeekCoder 33b, and Llama3.2 show limited effectiveness, fixing only a small number of vulnerabilities. The poor performance of these models could be attributed to several factors, such as insufficient exposure to Solidity code during pre-training or fine-tuning or architectures ill-suited for vulnerability repair, which requires a deep understanding of code syntax and security principles. The exceptionally poor performance of models like starcoder2 (marked with an asterisk *), along with incomplete data for opencoder, suggests potential issues with their training data or a fundamental mismatch between their capabilities and the task’s demands. These models might have been trained on an older version of Solidity or different smart contract security practices than those in the Not-So-Smart-Contracts dataset. Moreover, they might prioritize other aspects of code generation, such as code completion, over security-specific tasks like vulnerability repair.
Table 6: Move Vulnerability Repair (Time in seconds).
Model
UR
IL
UB
UC
UPF
UTC
Ov
PL
Time
CodeGeex4 6 0 0 6 1 1 1 0 96
CodeGemma 6 0 10 7 1 1 0 0 97
CodeLlama 15 0 1 17 2 2 2 1 244
CodeQwen 10 0 1 10 1 1 1 1 141
CodeStral 19 0 1 21 3 3 2 1 295
DeepSeekCoder 33b 26 0 2 30 4 4 3 1 412
DeepSeekV2 2 0 0 2 0 1 0 0 19
Gemma2 9b 7 0 1 8 1 1 0 10 108
Gemma2 27b 21 0 2 23 3 3 21 11 304
Granite-Code 6 0 0 6 1 1 1 0 90
Llama3.2 3 0 0 3 1 1 0 0 37
Llama3.3 70b 34 0 21 39 5 5 14 31 741
OpenCoder 7 0 1 8 1 1 0 0 94*
Qwen 2.5 code 5 0 0 6 1 1 1 0 80
StarCoder2 6 0 1 7 1 1 0 0 89
Smartify (Gemma2+CodeGemma) 293 2 16 189 41 48 51 10 112
Smartify (Gemma2+Llama3.1) 97 0 1 90 13 34 12 10 268
Move Prover dill2022fast - 2 - - - - 47 15 -
MoveLint - - - - 19 30 0 0 -
MoveScan song2024empirical 406 2 28 404 52 62 60 15 -
Abbreviations: UR: Unchecked Return; IL: Infinite Loop; UB: Unnecessary Boolean; UC: Unused Constant; UPF: Unused Private Function; UTC: Unnecessary Type Conversion; Ov: Overflow; PL: Precision Loss.
The public availability of the "Not So Smart Contracts" dataset raises the question of data contamination. Many evaluated models, especially those trained on large, public code corpora, might have encountered this dataset during pre-training, potentially inflating their performance. However, since CodeGemma and Smartify (Gemma2+CodeGemma) were specifically fine-tuned for this task, the issue of data contamination is likely less significant.
5.2 Move Code Repair
This section analyzes the efficacy of various models in repairing vulnerabilities within Move smart contracts, as detailed in Table 6. The evaluation encompasses eight distinct vulnerability categories: Unchecked Return (UR), Infinite Loop (IL), Unnecessary Boolean (UB), Unused Constant (UC), Unused Private Function (UPF), Unnecessary Type Conversion (UTC), Overflow (Ov), and Precision Loss (PL) following the works of Song et al song2024empirical. The metrics presented in the table represent the number of successfully repaired instances for each vulnerability type, with higher values indicating superior performance. The inference time, measured in seconds, is also provided for each model.
The results demonstrate a significant variance in performance across the evaluated models. Notably, the larger language models, such as Deepseekcoder 33b and Llama3.3 70b, exhibit a relatively higher number of successful repairs across multiple categories, albeit with a corresponding increase in inference time. Conversely, smaller models like DeepseekV2 and Llama3.2 demonstrate limited repair capabilities. The specialized tools for Move code, namely Move Prover, MoveLint, and MoveScan, were employed as a benchmark for comparison. It is crucial to note that these tools are designed for vulnerability detection rather than repair. MoveScan, in particular, identified a substantial number of instances across all categories, highlighting its effectiveness as a static analysis tool. Move Prover demonstrated proficiency in detecting Overflow and Precision Loss vulnerabilities, while MoveLint focused on Unused Private Functions and Unnecessary Type Conversions.
The Smartify models, which take advantage of a combination of Gemma2 with either CodeGemma or Llama3.1, present an interesting case. Smartify(Gemma2+CodeGemma) and Smartify(Gemma2+Llama3.1) outperform several individual models in multiple categories. This is likely because the specialized models are fine-tuned on the Move-specific dataset. For instance, Smartify (Gemma2+CodeGemma) achieves the highest number of repairs for the Unchecked Return, Infinite Loop, Unused Boolean, Unused Constant, Unused Private Function, Unnecessary Type Conversion, and Overflow categories, showcasing a substantial improvement over individual models in these areas. However, it is worth mentioning that they also have limitations compared to individual models for certain categories like Precision Loss.
RQ1 & RQ2 - Code Understanding and Vuln. Detection Yes. Our empirical analysis with Smartify, especially with using a fine-tuned Code-Gemma and also using vanilla pre-trained Llama3.1, has shown us the effectiveness of the framework’s ability to understand code and capture bad practices leading to vulnerability. Especially for a low-resource code like Move, without significant fine-tuning (in the case of Llama3.1), Smartify outperform the larger and computationally intensive models such as Llama 3.3 70b.
Notably, Smartify (Gemma2+CodeGemma), combining fine-tuned Gemma2 with CodeGemma, achieves performance on par with the best individual model, CodeGemma, which is expected due to one of the models being fine-tuned. This highlights the advantages of strategically combining specialized models, answering our next research question.
RQ3 & RQ4 - Code Repair Both for Solidity and Move, we were able to compare the efficacy of our framework with prior works. We can see Smartify outperforms all of the existing code models, even very specialized code models trained on Move (OpenCoder huang2024opencoder) in generating repair codes for detected vulnerabilities.
Furthermore, Smartify’s efficacy extends even when integrating a non-finetuned model like Llama 3.1. Smartify significantly outperforms Llama 3.2 by fixing 14 vulnerabilities compared to Llama 3.2’s single fix, making its performance comparable with the much more extensive and computationally intensive Llama 3.3 70b. This demonstrates that Smartify’s architecture can enhance even general-purpose language models for code repair, balancing speed, and accuracy.
RQ5 - Generalization Our implementation of Smartify with both fine-tuned Code-Gemma and Llama3.1 as the second agent allowed us to run our experiments on both sets of LLMs. The results show that Smartify can significantly boost performance even on non-finetuned models compared to a single model.
Comparative analysis reveals trade-offs between model scale and performance in automated code repair. Larger models, such as DeepSeekCoder 33b and Llama 3.3 70b, exhibit broader repair capabilities but incur higher computational costs and inference times. Conversely, the Gemma2 27b model demonstrates notable proficiency in addressing Overflow vulnerabilities, albeit with limitations in handling Unnecessary Boolean and Unused Constant compared to Llama 3.3 70b. While Llama 3.3 70b outperforms Smartify in overall repair capability, its significantly slower inference speed poses a challenge for practical deployment. Therefore, for real-world, on-device applications, Smartify (Gemma2+CodeGemma) presents a compelling solution with its balance of substantial accuracy and rapid inference.
Insight: Specialized code models like StarCoder lozhkov2024starcoder, OpenCoder huang2024opencoder and DeepSeekCoder guo2024deepseek doesn’t necessarily work well even if it’s a coding specific task. While code models like CodeGemma team2024codegemma and CodeLlama roziere2023code are much better at understanding instructions and working on code. This helped Smartify for its understanding and fine-tuning for code repairability.
Specialized static analysis tools for Move, including Move Prover, MoveLint, and MoveScan, work as baselines of detecting Move vulnerabilities with which we compare our Smartify and other LLMs. These findings underscore the need for targeted model improvements. The Smartify framework directly addresses these deficiencies, offering enhanced vulnerability repair effectiveness.
This research also opens up future research directions on using this framework for context-aware test case generation.
6 Ablation Study
To rigorously validate our architectural design and isolate the contribution of each key component, a comprehensive ablation study was conducted. This study systematically deconstructs the Smartify framework to quantify the impact of its core mechanisms—agent specialization, iterative refinement, and retrieval-augmented generation—on overall repair performance huang2025opencoderopencookbooktoptier; jiang2018contractfuzzer.
6.1 Ablation Configurations
Four distinct configurations of the system were evaluated against the Solidity and Move datasets:
  • Full Smartify: The complete five-agent system, including the RAG module for the Code Generator and the Refiner-Validator iterative feedback loop. This represents our proposed approach.
  • Smartify (No Refinement): A version of the framework where the Refiner and Validator agents are disabled. The output is taken directly from the first pass of the Code Generator. This configuration measures the impact of the iterative self-improvement loop on patch quality.
  • Smartify (No RAG): In this setup, the Code Generator operates without the contextual, few-shot examples provided by the RAG system. It relies solely on its fine-tuned knowledge and the Architect’s plan. This configuration is designed to measure the importance of providing language-specific, in-context examples, especially for the low-resource Move language.
  • Single-Agent Baseline (CoT-style): This configuration replaces the multi-agent system with a single, powerful agent (Gemma2 9B). The agent is given a complex, chained prompt that instructs it to sequentially perform the tasks of auditing, planning, and generating the repair. This mimics the linear reasoning workflow of CoT-based approaches like ContractTinker and directly tests the value of our role-based specialization and delegation architecture.
6.2 Analysis
The performance of each configuration was measured using the Pass@1 and Exploit Mitigation Rate metrics. The results, presented in Table 7, provide clear empirical evidence supporting our architectural choices.
Table 7: Ablation Study of Smartify Components on Solidity and Move Datasets
Solidity Move
Configuration Pass@1 (%) Exploit Mit. (%) Pass@1 (%) Exploit Mit. (%)
Full Smartify 26.7 25.0 48.9 45.7
Smartify (No Refinement) 25.0 20.0 44.6 39.1
Smartify (No RAG) 21.7 18.3 28.3 23.9
Single-Agent Baseline 18.3 15.0 21.7 17.4
The results demonstrate a clear and consistent degradation in performance as key components are removed, confirming the positive contribution of each element of the Smartify architecture.
  • Impact of Iterative Refinement: Removing the refinement loop (No Refinement) causes a noticeable drop in both metrics, particularly the Exploit Mitigation Rate (from 25.0% to 20.0% for Solidity). This indicates that while the initial code generation is often correct, the refinement process is crucial for hardening the patch against exploits and catching subtle regressions that the first pass might miss.
  • Impact of RAG: The removal of the RAG module (No RAG) has a significant negative impact across the board, but its effect is most pronounced for the Move language. The Pass@1 score for Move plummets from 48.9% to 28.3%, a relative decrease of over 42%. This strongly supports our hypothesis that providing in-context, language-specific examples is critical for achieving high performance in low-resource languages where the model’s pre-trained knowledge is limited.
  • Impact of Multi-Agent Architecture: The Single-Agent Baseline, designed to mimic a CoT-style approach, performs the worst of all configurations. Its performance is substantially lower than the Full Smartify framework, particularly on the more complex task of exploit mitigation. This finding provides strong evidence that our role-based, delegative architecture is superior to a monolithic, linear reasoning process. By specializing agents for distinct tasks (analysis vs. planning vs. generation), the system achieves a more robust and effective decomposition of the complex program repair problem.
The ablation study empirically validates the design of Smartify. The multi-agent architecture provides a superior structure for complex problem-solving, the refinement loop is essential for ensuring patch quality and security, and the RAG mechanism is a critical component for adapting the framework to new or low-resource programming languages.
7 Conclusion
This work addresses the pressing need for enhanced security in the burgeoning blockchain ecosystem. We investigate the application of Large Language Models (LLMs) to smart contract vulnerability detection and repair, focusing on Solidity and Move. We introduce Smartify, a novel multi-agent framework that significantly improves LLM performance in this critical domain. The contributions of this work are: (1) Smartify, a novel multi-agent framework that enhances LLM-based smart contract vulnerability detection and repair; (2) a method for encoding language-specific knowledge, valuable for low-resource languages like Move; (3) a scalable, adaptable approach applicable to other programming languages and LLMs; (4) a demonstration of Smartify’s efficacy on generalized pre-trained LLMs; and (5) a detailed analysis of the challenges inherent in automated code repair.
Smartify represents a significant advancement in automating smart contract security, a crucial concern in the expanding blockchain landscape. Future work will refine the framework, expand its language coverage, particularly within the blockchain domain, and integrate it into real-world blockchain development workflows. This research lays the foundation for AI-powered tools that can bolster the security and reliability of decentralized applications, fostering a more robust and trustworthy blockchain ecosystem.
Acknowledgments
This work was partially supported by the Sui Academic Research Award. We gratefully acknowledge the Google Developer Expert AI/ML team for providing cloud compute resources that facilitated this research.
We thank Sam Blackshear of Mysten Labs for his valuable feedback on the idea and early drafts of this paper.